OpenStudy (anonymous):

The number of real roots of (x+3)^4 + (x+5)^4 = 16 is: a) 0 b) 2 c) 4 d) none of these

OpenStudy (anonymous):

2

OpenStudy (king):

hw?

OpenStudy (anonymous):

homework

OpenStudy (king):

how?

OpenStudy (anonymous):

-3 and -5 I did inspection that's all I did, but you should expand it if you can

OpenStudy (king):

what do u mean by u did inspection?

OpenStudy (anonymous):

you can't do it, not now, maybe when you will have enough experience then... but for now just expand it

OpenStudy (anonymous):

(x + 3)^4 = 4^2 - ((x +5)^2)^2 now try it

OpenStudy (king):

ok

OpenStudy (anonymous):

That's rite, even I did it by seeing it in first glimpse. It is a DPP question of Vidyamandir Classes, NewDelhi. Is der any formula for finding the direct roots of a quartic equation?

OpenStudy (aravindg):

hmm surprising i think i saw this qn before wait,,

OpenStudy (anonymous):

Really? Then please answer something. Please.

OpenStudy (aravindg):

w8

OpenStudy (anonymous):

Okay bhai!!!!!!

OpenStudy (aravindg):

hey i got it aadarsh!!!

OpenStudy (anonymous):

I had posted the same question in maths group also. Got the answers.

OpenStudy (king):

the formula fr findin roots of a quadratic equation is \[\frac\]

OpenStudy (king):

\[\frac{-b \ \pm \sqrt{b ^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

\[(x+3)^4 + (x + 5)^4 = 16 \]\[x + 3 = t \tag1\]From (1), \[t^4 + (t+2)^4 = 16\]\[t^4 + (t+2)^4 = 2^4\]One can easily make out the equation is valid only when t is either 0 or -2.\[t = -2,0\]From (1), \(x = -5,-3\).